Help please

Help please
media%2Fd73%2Fd7322cf7-53c9-4a69-a0a3-bc
Answer #1

n = 100

p = 0.23

\mu = np = 100 * 0.23 = 23

\sigma = sqrt(np(1 - p))

     = sqrt(100 * 0.23 * 0.77)

    = 4.2083

P(X < 20)

= P(X < 19)

= P(X < 19.5)

= P((X - \mu)/\sigma< (19.5 - \mu)/\sigma)

= P(Z < (19.5 - 23)/4.2083)

= P(Z < -0.83)

= 0.2033

Option - A is correct.